OpenStudy (anonymous):

What is the slope, and what is the y-intercept of the line? 8x+6y=12

OpenStudy (anonymous):

Didn't I just tell you this? Again y = mx+ b, where m is your slope and b is the y - intercept, assuming you're asking a similar question. So solve for y.

OpenStudy (ilovecake):

This relationship always holds true: in y = mx + b, "b" gives the y-intercept, and "m" is the slope. We can use this fact to easily graph straight lines: Graph the equation y = ( 3/5 ) x – 2 from the slope and y-intercept.

OpenStudy (bee_see):